# for Hand Motion Generation in Piano Performance

### v1

@@ Appendix E Ethical Considerations
Piano motion datasets may pose significant privacy challenges, particularly concerning the pianist’s identifiable aspect, mainly their hands, during piano performance. Our dataset comprises videos uploaded by users on Bilibili, which are publicly accessible. To address privacy risks, we adopt a strict policy of releasing only video IDs, not the videos or images directly. Furthermore, we employ the MANO hand prior model [47] as a form of robust anonymization. This approach ensures the protection of personal information, mitigates the risk of individual identification, and minimizes privacy concerns. With these precautions, our dataset enables valuable research in piano motion analysis and generation.
@@ Appendix F Author Statement
Appendix G License and Consent with Public Resources
G.1 Tools for Annotation
The piano audios were transcribed by piano_transcription_inference [28] and the hand poses of the MANO model [47] in videos were annotated with MediaPipe [35] and HaMeR [42]:
G.2 Models for Baseline
Pre-trained Wav2Vec2.0 [4] and HuBERT [22] were utilized as audio feature extractor, while transformer [63] approach and SSM-based model [18] were employed as our feature decoder:
G.3 Re-evaluated Methods
In the experimental section, we evaluated EmoTalk [44] and LivelySpeaker [76], and models were reproduced using the official code:

### v2

@@ Appendix E Ethical Considerations
Piano motion datasets may pose significant privacy challenges, particularly concerning the pianist’s identifiable aspect, mainly their hands, during piano performance. Our dataset comprises videos uploaded by users on Bilibili, which are publicly accessible. During the dataset collection process, we carefully selected the high-quality piano performance videos with a bird’s-eye view on Bilibili and obtained permissions from the respective creators to use their videos, and accessed the original content. Those videos from creators who did not grant permission or did not respond have not been included in the dataset. We informed the creators that we would extract hand poses and audios from the videos to create a publicly available, non-commercial dataset. We clarified that their personal likenesses would not be used. The dataset also included metadata, such as the video URL, to trace the original video. In the camera-ready version, we will acknowledge all the creators’ contribution in building our dataset.